the concentration of FAS is 50.00mg/L and it is diluted to the 50.00mL mark on a volumetric flask what is the concentration of Fe(+2) in standard solution in mg/L? and what is the concentration for solutions with 2mL, 5mL, and 10mL (all using a 50.00mL flask)?

-The original concentration for 50 mg/L is 50.00mL as indicated in bold above
Answer #1

1. 2 ml

mass of FAS in this sample = concentration * volume of sample = 50 mg/L * 2 / 1000 = 0.1 mg

After diluting to 50 ml, concentration = mass of FAS / Volume of FAS = 0.1 mg / (50 * 10-3 L) = 2 mg/L

2. 5 ml

mass of FAS in this sample = concentration * volume of sample = 50 mg/L * 5 / 1000 = 0.25 mg

After diluting to 50 ml, concentration = mass of FAS / Volume of FAS = 0.25 mg / (50 * 10-3 L) = 5 mg/L

3. 10 ml

mass of FAS in this sample = concentration * volume of sample = 50 mg/L * 10 / 1000 = 0.5 mg

After diluting to 50 ml, concentration = mass of FAS / Volume of FAS = 0.5 mg / (50 * 10-3 L) = 10 mg/L